OpenStudy (anonymous):

Linear Algebra Question: Find a basis for nullspace(A) A = | 2 6 -3 1 | |-2 -3 2 0 | |-2 -3 2 0 | | 0 3 -1 1 | The answer I get is [ 1 3 (-3/2) (1/2) ] , [ 0 1 (-1/3) (1/3) ] and it is wrong. What do you guys get?

OpenStudy (zarkon):

\[\left[\begin{matrix}1/2\\-1/3\\0 \\ 1\end{matrix}\right],\left[\begin{matrix}1/2\\1/3\\1 \\ 0\end{matrix}\right]\]

OpenStudy (anonymous):

how? :P

OpenStudy (anonymous):

nevermind; i figured it out :D thanks